calculate the number of moles in each of the following 1) 12.6g water 2) 12.6g sodium chloride 3) 4.57x10^17
1
Expert's answer
2012-07-26T12:00:43-0400
1)
The molar mass of water is 18 g/mol
The amount of water is n(H2O) =
m(H2O)/M(H2O) = 12.6 g / 18 g/mol = 0.7 mol
2)
The molar mass of NaCl is
58.5 g/mol
The amount of water is n(NaCl) = m(NaCl)/M(NaCl) = 12.6 g / 58.5
g/mol = 0.215 mol
3)
n(particles) = N(particles) / N(Avogadro) = 4.57*1^17
/ 6.022*10^23 = 7.6*10^-7 mol